Lars:

The volume of a cube can be found using the equation V = s³, where V is the volume and s is the measure of one side of the cube. Match the equation for how to solve for the side length of a cube to its description. Drag the equation into the box to match the description.

Lars:

Luigi0210:

Since the equation is \[\Large V=S^3 \] in order to find the side "S" you have to take the CUBIC root of both sides \[\Large \sqrt[3]{V} =S \] Since V = 756 \[\Large \sqrt[3]{756} = S\]

Lars:

@luigi0210 wrote:
Since the equation is \[\Large V=S^3 \] in order to find the side "S" you have to take the CUBIC root of both sides \[\Large \sqrt[3]{V} =S \] Since V = 756 \[\Large \sqrt[3]{756} = S\]
Thank you.